OpenStudy (anonymous):

Solve the proportion. 2/x=sqrt3+1/5

OpenStudy (saifoo.khan):

\[\frac 2x = \sqrt 3 + \frac 15\] Like this?

OpenStudy (anonymous):

no hold on ill show you

OpenStudy (anonymous):

no hold on ill show you\[\frac{ 2 }{ x }=\frac{ \sqrt3+1 }{ 5 }\]

OpenStudy (anonymous):

 why was that so easy imust be wrong somewhere or u must be kidding me

OpenStudy (anonymous):

@jess.white

OpenStudy (anonymous):

no that was it

OpenStudy (anonymous):

Thank you though maybe you can help on my next one

OpenStudy (anonymous):

was that the answer

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

ur welcome